# A simple, flexible method for timing cross-calibration of space missions

### v1

@@ Abstract
The timing (cross-)calibration of astronomical instruments is often done by comparing pulsar times-of-arrival (TOAs) to a reference timing model. In high-energy astronomy, the choice of solar system ephemerides and source positions used to barycenter the photon arrival times has a significant impact on the procedure, requiring a full reprocessing the data each time a new convention is used. Our method, developed as part of the activities of the International Astronomical Consortium for High Energy Calibration (IACHEC), adapts an existing pulsar solution to arbitrary JPL ephemerides and source positions by simulating geocentric TOAs and refitting timing models (implemented with PINT). We validate the procedure and apply it to thousands of observations of the Crab pulsar from 14 missions spanning 2002–2025, demonstrating inter-ephemeris TOA consistency at the $\lesssim 5\rm\mu s$ level, using the DE200/FK5-based Jodrell Bank Monthly Ephemeris as a reference. We release open-source tools (TOAextractor) and a TOA database to support future calibration and scientific studies. Instrument timing performance is broadly consistent with mission specifications; the X-ray-to-radio phase offset varies with energy and time at a level that is marginally compatible with the uncertainties of the radio ephemeris, motivating coordinated multiwavelength follow-up.
@@
facilities: CXO, Einstein Probe, Fermi, Hitomi, HXMT, IXPE, MAXI, NICER, Ninjasat, NuSTAR, Suzaku, Swift(XRT and BAT), XMM-Newton, XRISMsoftware: astropy (Astropy Collaboration et al., 2013, 2018, 2022), Stingray (D. Huppenkothen et al., 2019), PINT (J. Luo et al., 2021), TOA extractor (M. Bachetti, 2025b)
1 Introduction
@@
The Timing Working Group (TWG) focuses on the temporal cross-calibration of different missions. The goal of the TWG is to develop methods to compare the timing performance of different missions, using pulsars as a common clock. Pulsars are neutron stars (NSs) that emit beams of radiation swiping the sky like a lighthouse (giving the characteristic pulsed signal). Discovered in the radio band (A. Hewish et al., 1968), pulsars have later been detected at all wavelengths, from radio to gamma-rays (R. B. E. Lovelace et al., 1968; H. Tananbaum et al., 1972; J. Nelson et al., 1970; S. K. Gupta et al., 1978; A. A. Abdo et al., 2013).
[…]
@@
NSs, however, are probably not homogeneous spheres, but they likely have a complex internal structure, with a solid crust and a liquid core. “Glitches” in their spin evolution, that produce sudden changes of spin period or its derivatives, are thought to arise exactly from the pinning and unpinning of the internal layers of the star (see R. N. Manchester, 2018, for a review). Accretion from a companion star can also change the spin period of a pulsar, either by adding angular momentum to the star (spin-up) or by removing it (spin-down) (P. Ghosh & F. K. Lamb, 1978). The fastest-rotating pulsars, known as millisecond radio pulsars (MSPs), are usually old pulsars that have gone through a re-acceleration process through accretion and have a low magnetic field, which limits the rotational noise and further stabilizes the rotation. In fact, some MSPs have rotations that can be predicted with a precision of microseconds in a few years (e.g. D. J. Reardon et al., 2024).
@@
Even though the precise emission mechanisms of pulsars are still debated, NSs have a very strong surface magnetic field (from 108 to $\sim$1015 G), which is responsible for most of their phenomenology. First of all, the energy loss from the rotating dipole field that slows pulsars down is also believed to power most radio pulsars (for this reason, also called rotation-powered pulsars, or RPPs) with $B\lesssim 10^{12}G$ (D. R. Lorimer, 2008). Accreting pulsars are instead powered by the accretion of matter which is funneled on the surface by the magnetic field (P. Ghosh & F. K. Lamb, 1978), producing a bright X-ray emission. The highest-magnetic-field pulsars, known as magnetars, are believed to be powered by the decay of the magnetic field itself, and they show a complicated phenomenology, including giant flares and bursts at high energies (N. Rea & D. De Grandis, 2025).
@@
For their high rotational stability, pulsars have often been used as standard clocks, with applications ranging from the detection of gravitational waves (J. H. Taylor & J. M. Weisberg, 1982; G. Agazie et al., 2023; EPTA Collaboration et al., 2023) to space navigation (K. D. Anderson et al., 2015; A. A. Emadzadeh & J. L. Speyer, 2011). Another application has been the timing calibration of astronomical instruments (e.g. G. Cusumano et al., 2012; A. Martin-Carrillo et al., 2012; J. S. Deneva et al., 2019; M. Bachetti et al., 2021).
@@
PSR B0531+21, also known as PSR J0534+2200 or the Crab pulsar due to its position at the center of the Crab nebula M1, is a RPP with a spin period of $P\sim 33$ ms, and a spin-down rate of $\dot{P}\sim 4.2\times 10^{-13}$ s s-1, which corresponds to a spin-down luminosity of $\dot{E}=4\pi^{2}I\dot{P}/P^{3}\sim 4\times 10^{38}$ erg s-1 (D. R. Lorimer, 2008). Despite being young [^1] The Crab nebula has been identified as the remnant of SN 1054, recorded in 1054 A.D. by astronomers all around the globe (N. U. Mayall, 1939) and not particularly stable by pulsar standards (including glitches and a variable spin down rate that amount to a large “timing noise”), it is one of the best studied RPPs in the sky, due to being very bright across the electromagnetic spectrum and, not less importantly, due to the Crab nebula’s relatively stable (and power law-like in the X-rays) spectrum which made it an excellent flux calibrator. This provides pulsar astronomers with a wealth of “free” observations of the pulsar. The pulsar’s spin evolution has been tracked for four decades by the Jodrell Bank Observatory (A. G. Lyne et al., 1993), which still today provides a monthly ephemeris (the Jodrell Bank Monthly Ephemeris, hereafter JBE) of the pulsar’s spin frequency.
@@
This makes the Crab pulsar an excellent tool for tracking the timing calibration of X-ray instruments, and it has been used as such by many missions (e.g. L. Kuiper et al., 2003; A. H. Rots et al., 2004; D. A. Smith et al., 2008; Y. Terada et al., 2008; S. Molkov et al., 2010; G. Cusumano et al., 2012; A. Martin-Carrillo et al., 2012; M. Bachetti et al., 2021; A. Basu et al., 2021; Y. Tuo et al., 2022; G. Cusumano et al., 2024).
@@
Typically, these missions use the JBE to fold the pulsar data and obtain pulse profiles and TOAs, which are then compared to the expected TOAs from the ephemerides. The main problem with this approach is that the Jodrell Bank ephemerides are based on the DE200 solar system ephemeris (E. M. Standish, 1982), which is quite outdated, and they do not account for proper motion, so assuming that the pulsar’s position is fixed in the sky. This makes sense to provide a consistent timing solution for the Crab pulsar, but it is not ideal for cross-calibration purposes, as it forces all missions to use the same ephemeris and source position, losing flexibility when gathering data from different missions, and often needing to perform separate data processing for timing calibration and scientific analysis.
@@
In this paper, we present a new method for comparing the times of arrival (TOAs) of pulsar signals obtained with different instruments, even when they use different JPL ephemerides, source positions, and reference frames. We apply the method to the temporal cross-calibration of different missions, using the Jodrell Bank Montly Ephemeris of the Crab pulsar as a common reference.
2 Processing
2.1 Typical pulsar timing workflow
@@
The pulsar signals are periodic in nature, repeating at the pulsar’s spin period, which is typically in the range of milliseconds to seconds (with a few cases of pulsars with spin periods up to hours, see the reviews D. R. Lorimer 2008; N. Rea & D. De Grandis 2025). It is common to express the signal at a time $t$ in terms of the pulsar’s spin phase $\phi(t)$, defined through the pulsar’s spin frequency $\nu$ and its time derivative $\dot{\nu}$, $\ddot{\nu}$ with the following Taylor expansion:
@@
The single pulses of pulsars are typically too faint to be detected directly, and they have highly variable shapes, so the data are folded at the pulsar’s spin period to produce an average pulse profile which is instead quite stable and constitutes a recognizable “fingerprint” for a given pulsar. Folding is done by “cutting” the time series in identical segments whose length is the pulsar’s spin period. The segments are then averaged together, and if the segments are of exactly the right length, the signal-to-noise ratio of the resulting profile is maximized. When the time series is constituted by photons, folding is done by doing a histogram of the phase of the photon, basically the fraction of the pulsar’s spin period that has elapsed since the last pulse.
@@
A new TOA is usually defined as the time of arrival of the peak of the folded pulse profile, and that gives phase 0. The maximum can be determined precisely by modeling the pulse profile with a series of Gaussian or equivalent curves. When calculating TOAs from a given instrument, if the pulse profile does not evolve with time, the procedure is generally done by calculating a cross-correlation with a template profile and/or using the FFTFIT algorithm that is a chi-square fit of the normalized template to the folded profile in the frequency domain (J. H. Taylor, 1992). The template is usually a model of a high signal-to-noise profile of the same pulsar.
2.2 X-ray and Gamma-ray data processing
@@
Pulsar timing requires to pay extreme attention to the detail of how the times of our measurements are recorded.
@@
This Paper uses data from 14 missions, each with its own specific procedure for data reduction, which usually follows the same pattern: using multiple sources of data, from “housekeeping” measurements recorded on the spacecraft to space weather information, mission-specific pipelines produce “cleaned” datasets in the form of “event files”, which are typically FITS files with a table containing a time stamp and a number of physical properties for each detection that fulfils some quality criteria. The details of each mission’s pre-processing pipelines can be found in the Appendix A.
@@
The time stamps of events are usually expressed in the mission elapsed time (MET), which is a topocentric time at the instrument location expressed in seconds from a reference modified julian date (MJD), called MJDREF. The MJDREF is written in the header of the FITS file, either as a single keyword, or as the combination of its integral part MJDREFI and the fractional part MJDREFF. Often, a TIMEZERO keyword adds a global correction, expressed in seconds, to the event times. In detectors where the measurement duration (hereafter “frame time”) is significant, the TIMEDEL keyword is used to indicate the frame time, and TIMEPIXR to specify if the timestamps are referred to the start (0), the middle (0.5) or the end (1) of the frame time. Finally, some missions carry time references whose alignment to UT drifts (e.g. quartz oscillators whose frequency can change depending on temperature), and for these missions an additional correction to their time reference can be provided on the ground. The total terrestrial time (TT) measured at the spacecraft can then be calculated as [^2] https://heasarc.gsfc.nasa.gov/docs/xte/abc/time_tutorial.html MJD(TT) = (MJDREFI + MJDREFF) + (TIME + TIMEZERO + CLOCKCORR + (0.5 - TIMEPIXR) * TIMEDEL) / 86400
Depending on the position of the Earth in its orbit and the position of instrument around the Earth, the arrival time at the instrument of a given photon can change by up to 16 minutes in a year and by milliseconds in an orbital timescale of the satellite, so it is important to find a way to express the time of arrival at a fixed reference point, typically the solar system barycenter (SSB).
@@
In radio pulsar studies, it is common to calculate the arrival time of pulsations through folding in the topocentric frame (so, as a UT time at the location of the radio telescope) using atomic clocks, and then to convert to the SSB during the modeling phase, depending on the version of the ephemeris used in the model. This is convenient because the position of the instrument is fixed with respect to the Earth, which rotates relatively slowly.
@@
In X-ray missions, the position of the instrument changes rapidly due to the orbit of the satellite, so it is not practical to use a topocentric time for the folding procedure (as it often requires to accumulate data over multiple orbits). Therefore, in X-ray astronomy the conversion of times to the SSB, or barycentering, often happens before the folding, using tools like the barycorr FTOOL together with information on the satellite’s position and velocity encoded in the orbit/attitude files shared by science operation centers as part of the “housekeeping” data for each observation.
@@
The coordinates of the source are a key parameter for barycentering, as we are going to detail in Section 2.3. Different barycentering tools encode the coordinates used in different FITS keywords (e.g. RA,DEC_OBJ for barycorr, RA,DEC_TDB for HEASOFT’s barycen).
@@
All the relevant data are retrieved from the FITS files using Stingray (D. Huppenkothen et al., 2019).
2.3 Need for time system consistency

### v3

@@ Abstract
The timing (cross-)calibration of astronomical instruments is often done by comparing pulsar times-of-arrival (TOAs) to a reference timing model. In high-energy astronomy, the choice of solar system ephemerides and source positions used to barycenter the photon arrival times has a significant impact on the procedure, requiring a full reprocessing of the data each time a new convention is used. Our method, developed as part of the activities of the International Astronomical Consortium for High Energy Calibration (IACHEC), adapts an existing pulsar solution to arbitrary JPL ephemerides and source positions by simulating geocentric TOAs and refitting timing models (implemented with PINT). We validate the procedure and apply it to thousands of observations of the Crab pulsar from 15 missions spanning 1996–2025, demonstrating inter-ephemeris TOA consistency at the $\lesssim 5\rm\mu s$ level, using the DE200/FK5-based Jodrell Bank Monthly Ephemeris as a common reference. We release the TOAExtractor open-source tool and a TOA database to support future calibration and scientific studies. Instrument timing performance is broadly consistent with mission specifications; the X-ray-to-radio phase offset varies with energy and time at a level that is marginally consistent with the uncertainties of the radio ephemeris, motivating coordinated multiwavelength follow-up.
@@
facilities: CXO, EP, Fermi, Hitomi, HXMT, IXPE, MAXI, NICER, Ninjasat, NuSTAR, Suzaku, Swift (XRT and BAT), XMM, XRISM, XTE.software: astropy (Astropy Collaboration et al., 2013, 2018, 2022), Stingray (Huppenkothen et al., 2019), PINT (Luo et al., 2021), TOA extractor (Bachetti, 2025c)
I Introduction
@@
The Timing Working Group (TWG) focuses on the temporal cross-calibration of different missions. The goal of the TWG is to develop methods to compare the timing performance of different missions, using pulsars as a common clock. Pulsars are neutron stars (NSs) that emit beams of radiation swiping the sky like a lighthouse (giving the characteristic pulsed signal). Discovered in the radio band (Hewish et al., 1968), pulsars have later been detected at all wavelengths, from radio to gamma-rays (Lovelace et al., 1968; Tananbaum et al., 1972; Nelson et al., 1970; Gupta et al., 1978; Abdo et al., 2013).
With the mass of 1-2 $M_{\odot}$ and a radius of about 10 km, NSs are the densest objects in the universe excluding stellar-mass black holes. Being so compact, they are effectively point-like as seen from nearby stars, and tidal forces are usually negligible. Their rotational speed can change over time due to a number of intrinsic and external factors. The emission of electromagnetic radiation and a particle wind from their rotating magnetic field produces a characteristic spin-down.
@@
NSs, however, are probably not homogeneous spheres, but they likely have a complex internal structure, with a solid crust and a superfluid core. “Glitches” in their spin evolution, that produce sudden changes of spin period or its derivatives, are thought to arise exactly from the pinning and unpinning of the internal layers of the star (see Manchester, 2018, for a review). Accretion from a companion star can also change the spin period of a pulsar, either by adding angular momentum to the star (spin-up) or by removing it (spin-down) (Ghosh and Lamb, 1978). The fastest-rotating pulsars, known as millisecond radio pulsars (MSPs), are usually old pulsars that have gone through a spin-up process through accretion and have a low magnetic field, which limits the rotational noise and further stabilizes the rotation. In fact, some MSPs have rotations that can be predicted with a precision of microseconds in a few years (e.g. Reardon et al., 2024).
@@
Even though the precise emission mechanisms of pulsars are still debated, NSs have a very strong surface magnetic field (from 108 to $\sim$1015 G), which is responsible for most of their phenomenology. First of all, the energy loss from the rotating dipole field that slows pulsars down is also believed to power most radio pulsars (for this reason, also called rotation-powered pulsars, or RPPs) with $B\lesssim 10^{12}$ G (Lorimer, 2008). Accreting pulsars are instead powered by the accretion of matter which is funneled on the surface by the magnetic field (Ghosh and Lamb, 1978), producing a bright X-ray emission. The highest-magnetic-field pulsars, known as magnetars, are believed to be powered by the decay of the magnetic field itself, and they show a complicated phenomenology, including giant flares and bursts at high energies (Rea and De Grandis, 2025).
@@
For their high rotational stability, pulsars have often been used as standard clocks, with applications ranging from the detection of gravitational waves (Taylor and Weisberg, 1982; Agazie et al., 2023; EPTA Collaboration et al., 2023) to space navigation (Sheikh et al., 2006; Emadzadeh and Speyer, 2011; Anderson et al., 2015; Mitchell et al., 2018; Yidi et al., 2023). Even more intriguing, pulsars have been proposed as reference clocks to establish a new time scale, independent of atomic clocks (e.g. Hobbs et al., 2020). Another application, which is the subject of this work, has been the timing calibration of astronomical instruments (e.g. Kuiper et al., 2003; Rots et al., 2004; Smith et al., 2008; Terada et al., 2008b; Molkov et al., 2010; Martin-Carrillo et al., 2012; Cusumano et al., 2012; Deneva et al., 2019; Bachetti et al., 2021; Basu et al., 2021; Tuo et al., 2022; Cusumano et al., 2024).
@@
PSR B0531+21, also known as PSR J0534+2200 or the Crab pulsar due to its position at the center of the Crab nebula M1, is a RPP with a spin period of $P\sim 33$ ms, and a spin-down rate of $\dot{P}\sim 4.2\times 10^{-13}$ s s-1, which corresponds to a spin-down luminosity of $\dot{E}=4\pi^{2}I\dot{P}/P^{3}\sim 4\times 10^{38}$ erg s-1 (Lorimer, 2008). Despite being young [^1] The Crab nebula has been identified as the remnant of SN 1054, recorded in 1054 A.D. by astronomers all around the globe (Mayall, 1939) and not particularly stable by pulsar standards (including glitches and a variable spin down rate that amount to a large “timing noise”), it is one of the best studied RPPs in the sky, due to being very bright across the electromagnetic spectrum and, not less importantly, due to the Crab nebula’s relatively stable (and power law-like in the X-rays) spectrum which made it an excellent flux calibrator. This provides pulsar astronomers with a wealth of “free” observations of the pulsar. The pulsar’s spin evolution has been tracked for four decades by the Jodrell Bank Observatory (Lyne et al., 1993), which still today provides a monthly ephemeris (the Jodrell Bank Monthly Ephemeris, hereafter JBE) of the pulsar’s spin frequency.
@@
This makes the Crab pulsar an excellent tool for tracking the timing calibration of X-ray instruments, and it has been used as such by many missions (e.g. Kuiper et al., 2003; Rots et al., 2004; Smith et al., 2008; Terada et al., 2008b; Molkov et al., 2010; Cusumano et al., 2012; Martin-Carrillo et al., 2012; Bachetti et al., 2021; Basu et al., 2021; Tuo et al., 2022; Cusumano et al., 2024).
@@
Typically, these missions use the JBE to fold the pulsar data and obtain pulse profiles and TOAs, which are then compared to the expected TOAs from the ephemerides. The main problem with this approach is that the Jodrell Bank ephemerides are based on the DE200 solar system ephemeris (Standish, 1982), which is quite outdated, and they do not account for proper motion, so assuming that the pulsar’s position is fixed in the sky. This makes sense to provide a consistent timing solution for the Crab pulsar, but it is not ideal for cross-calibration purposes, as it forces all missions to use the same ephemeris and source position, losing flexibility when gathering data from different missions, and often needing to perform separate data processing for timing calibration and scientific analysis.
@@
In this paper, we present a new method for comparing the times of arrival (TOAs) of pulsar signals obtained with different instruments, even when they use different JPL ephemerides, source positions, and reference frames. We apply the method to the temporal cross-calibration of different missions, using the Jodrell Bank Monthly Ephemeris of the Crab pulsar as a common reference. In section II, we describe the typical pulsar timing workflow and the specific data processing steps for X-ray and gamma-ray data. In section III, we describe our method for adapting pulsar timing solutions to different ephemerides and source positions. In section IV, we present the results of applying our method to a large dataset of Crab pulsar observations from multiple missions. Finally, in section V, we summarize our findings and discuss the implications for future timing calibration efforts. Appendix A contains definitions of time scales and conventions used in pulsar timing, Appendix B contains details on the data processing for each mission used in this work.
II Processing
II.1 Typical pulsar timing workflow
@@
The pulsar signals are periodic in nature, repeating at the pulsar’s spin period, which is typically in the range of milliseconds to seconds (with a few cases of pulsars with spin periods up to hours, see the reviews Lorimer 2008; Rea and De Grandis 2025). It is common to express the signal at a time $t$ in terms of the pulsar’s spin phase [^2] We express the phase from 0 to 1. Other works might use 0-$2\pi$. $\phi(t)$, defined through the pulsar’s spin frequency $\nu$ and its time derivative $\dot{\nu}$, $\ddot{\nu}$ with the following Taylor expansion:
@@
The single pulses of pulsars are typically too faint to be detected directly, and, especially in the radio band, they can have highly variable shapes, so the data are folded at the pulsar’s spin period to produce an average pulse profile which is instead quite stable and constitutes a recognizable “fingerprint” for a given pulsar. Folding is done by “cutting” the time series in identical segments whose length is the pulsar’s spin period. The segments are then averaged together, and if the segments are of exactly the right length, the signal-to-noise ratio of the resulting profile is maximized. When the time series is constituted by photons, folding is done by doing a histogram of the phase of the photon, basically the fraction of the pulsar’s spin period that has elapsed since the last pulse.
@@
A new TOA is usually defined as the time of arrival of the peak (or another fiducial point) of the folded pulse profile, and that gives phase 0. The maximum can be determined precisely by modeling the pulse profile with a series of Gaussian components or equivalent curves. When calculating TOAs from a given instrument, if the pulse profile does not evolve with time, the procedure is generally done by calculating a cross-correlation with a template profile and/or using the FFTFIT algorithm that is a chi-square fit of a linear phase gradient between the Fourier representations of a template and the folded profile (Taylor, 1992). The template is usually a model of a high signal-to-noise profile of the same pulsar. In high-energy astronomy, alternative likelihood-based approaches are sometimes used (Livingstone et al., 2009; Ray et al., 2011; Clark et al., 2015; Nieder et al., 2019, e.g.). In this paper, we use a method based on maximum-likelihood fitting of pulse profiles (avoiding the a priori definition of a template), see Section III.3.
II.2 X-ray and Gamma-ray data processing
@@
Pulsar timing requires extreme attention to the detail of how the measurement times are recorded.
@@
This paper uses data from 15 missions, each with its own specific procedure for data reduction, which usually follows the same pattern: using multiple sources of data, from “housekeeping” measurements recorded on the spacecraft to space weather information, mission-specific pipelines produce “cleaned” datasets in the form of “event files”, which are typically FITS files with a table containing a time stamp and a number of physical properties for each detection that fulfils some quality criteria. The details of each mission’s pre-processing pipelines can be found in the Appendix B.
@@
The time stamps of events are usually expressed in the mission elapsed time (MET), which give the number of seconds from a reference modified julian date (MJD), called MJDREF, typically using the Terrestrial Time scale (for definitions of time scales and conventions, see Appendix A). The MJDREF is written in the header of the FITS file, either as a single keyword, or as the combination of its integral part MJDREFI and the fractional part MJDREFF [^3] https://heasarc.gsfc.nasa.gov/docs/heasarc/ofwg/docs/summary/ogip_93_003_summary.html. When this happens, sometimes the MJDREFF encodes the number of leap seconds (divided by 86400 to express them in days). In this case, using just MJDREFI can give times in the UTC scale, while summing MJDREFF + MJDREFI gives times in the TT scale. Often, a TIMEZERO keyword adds a global correction, expressed in seconds, to the event times. In detectors where the measurement duration (hereafter “frame time”) is significant, the TIMEDEL keyword is used to indicate the frame time, and TIMEPIXR to specify if the timestamps are referred to the start (0), the middle (0.5) or the end (1) of the frame time. Finally, some missions carry time references whose alignment to UT drifts (e.g. quartz oscillators whose frequency can change depending on temperature), and for these missions an additional correction to their time reference can be provided on the ground. We refer to this additional term as FINECLOCK below. The total time in the TT scale, measured at the spacecraft, can then be calculated as [^4] https://heasarc.gsfc.nasa.gov/docs/xte/abc/time_tutorial.html MJD(TT) = (MJDREFI + MJDREFF) + (TIME + TIMEZERO + FINECLOCK + (0.5 - TIMEPIXR) * TIMEDEL) / 86400. Different time scales are allowed by the FITS standard, provided that the TIMESYS keyword is different from TT. As mentioned before, in some missions MJD(UTC) can be calculated by just omitting the MJDREFF value in the formula above.
@@
In radio pulsar studies, it is common to calculate the arrival time of pulsations through folding in the topocentric frame (so, as a UT time at the location of the radio telescope) using atomic clocks as a reference, and then to convert to the SSB during the modeling phase, depending on the version of the ephemeris used in the model.
@@
In high-energy missions, the detection of pulses often requires integration over multiple orbits of the satellites, making topocentric TOAs impractical. Therefore, in X-ray astronomy the conversion of times to the SSB, or barycentering, often happens before the folding, using tools like the barycorr FTOOL together with information on the satellite’s position and velocity encoded in the orbit/attitude files shared by science operation centers as part of the “housekeeping” data for each observation.
@@
The coordinates of the source are a key parameter for barycentering, as we are going to detail in Section II.3. Different barycentering tools encode the coordinates used in different FITS keywords (e.g. RA,DEC_OBJ for barycorr, RA,DEC_TDB for HEASOFT’s barycen).
@@
All the relevant data are retrieved from the FITS files using Stingray (Huppenkothen et al., 2019).
II.3 Consistency in ephemerides, coordinates, reference frames